OpenStudy (anonymous):

Graph the given function. Identify the vertex and the axis of symmetry y=2x^2+8x+9 I ended up with a vertex of (-2,7) but it should be (-2,1). Can somone show me how to get that?

OpenStudy (campbell_st):

well the easy way to find the vertex is to 1. find the line of symmetry \[x = \frac{-b}{2a}\] you have b = 8 and a = 2 so \[x = \frac{-8}{ 2 \times 2}\] then substitute the answer into the equation... so be careful with the substitution \[y = 2 \times (-2)^2 + 8 \times (-2) + 9\] hope it helps

OpenStudy (anonymous):

Thank you! :)